How do I do this? 1. When a particular baseball players come to the plate, he has a % chance of getting a hit. During one particular game, the
Answer #1

Let X be the number of hit in his 3 bats. Then X ~ Binomial(n = 3, p = 1/3)

A.

Probability that the player gets exactly 1 hit in 3 bats = P(X = 1) = 3C1 * (1/3)1 * (1 - 1/3)3-1

= 3 * (1/3) * (2/3)2

= 0.4444

B.

Probability that the player gets at least 1 hit = P(X \ge 1) = 1 - P(X = 0)

= 1 - 3C0 * (1/3)0 * (1 - 1/3)3-0

= 1 - (2/3)3

= 0.7037

C.

Probability that the player has two hits in his first two at bats = (1/3) * (1/3) = 1/9